Factor the expressions completely.
x^2-9
x^2+6x+9
5x^2+16x+3

Answer by ewatrrr(24785)

 
Hi,
x^2-9 = (x-3)(x+3)
x^2+6x+9 = (x+3)(x+3)
5x^2+16x+3 = (5x +1)(x+3)
Check the factoring for correctness by using FOIL
F First terms
O Outside terms
I Inside terms
L Last terms